Suppose the first-order maximum for 475 nm blue light being transmitted through a diffraction grating is at an angle of 22°.

Randomized Variables
λ = 475 nm
θ = 22°


How many lines per centimeter are there on this diffraction grating?
1
Expert's answer
2020-12-02T07:33:09-0500

Answer

Equation of mth order maximum

"d\sin \theta =m\lambda"


"d =\frac{m\lambda}{\sin \theta}=\frac{1\times475\times10^{-9}}{\sin 22^°}"

"d=1268\times10^{-9}m"

lines per centimeter is inverse of spacing

"N=\frac{1}{1268\times10^{-9}m}=7886" lines